# Density of wave fronts

### v1

@@ 1. Motivating questions
These pictures illustrate some rather general wave front problems in Riemannian geometry: for which Riemannian manifolds with or without boundary or for which polytopes do the wave front become dense for all initial points, like in the mirrored room example? In which cases does the wave front become dense for some points and not become dense for others like in the duck example case? In which cases does it not become dense for all initial points, like in the tsunami case?
@@
We explore this theme here in the torus, square and cube surface case, where the metric is flat and where the question can be analyzed with elementary tools. While the subject could reach far into differential geometry, all we really need here is some single variable calculus.
Figure 1. What happens with a light wave front in a room of mirrors, a tsunami on a perfect sphere or a cube, and the wave front produced by a duck in a circular pond.
@@ 2. Background
The third framework considers piecewise smooth Riemannian polytops, a concept historically challenging to define [30, 26]. This difficulty has led either to restricting to convex polytopes [19, 38] or to use a broader definition, where a polytop $M$ is viewed as the geometric realization of a finite abstract simplicial complex $G$, a finite set of non-empty sets closed under the operation of taking non-empty subsets [17]. A piecewise smooth Riemannian polytop is such a polytop that is equipped with a Riemannian metric $g$ that is smooth except on a subset of codimension $2$. Inductively, this means $M$ is a smooth manifold (with or without boundary) for which the tensor $g$ is defined on an open subset $U=M\setminus K$, where $K$ is a piecewise smooth Riemannian polytope of codimension $2$. The geodesic flow is defined for geodesic paths avoiding $K$. One of the simplest examples is the square table, a $2$-manifold with circular boundary and a metric discontinuous at its four corners. For a $3$-dimensional example, consider a billiard in a solid cube. Here, the Riemannian metric tensor $g$ may be discontinuous along the edges and vertices forming a 1-dimensional network.
@@ 3. Wave fronts
  • 2. convex planar flat billiards. Now, $W_{t}$ remains a continuous curve. It remains piecewise smooth. Differentiability can fail at caustics or the boundary
  • 3. piecewise flat convex polyhedra In this case, $W_{t}$ in general can already become disconnected.
Figure 2. In a Riemannian d-manifold $M$, the wave front $W_{t}(P)$ of a point $P$ for which a tangent space $T_{P}M$ exists, is the image $\exp_{P}(S_{t})$ of the sphere $S_{t}=\{|x|=t\}\subset T_{P}M\sim\mathbb{R}^{d}$ of radius $t$.
@@
While for a smooth surface $M$ without boundary or a billiard in a convex region, the wave fronts $W_{t}$ are piecewise smooth and in particular continuous, this is already different on a closed polyhedral surface. The wave front there can become disconnected after some time. This happens in particular for the cube surface, the surface bounding the unit cube. Numerical experiments still indicate, also for more general non-flat compact surfaces, that in general, the wave front becomes dense, meaning that it fills out the surface more and more.
3.3.
Here is the definition of ”dense wave front” for a general compact Riemannian manifold $M$ with or without boundary. Given $r>0$, the open ball $B_{r}(Q)$ of a point $Q\in M$ is the set of points $P$ for which there is a geodesic from $P$ to $Q$ that has length smaller than $r$.
@@
Since the $(d-1)$-dimensional volume of the co-dimension-one wave front $W_{t}$ is expected to grow and the manifold $M$ is finite, we expect it to fill up $M$ more and more. The term “generic” is here used in an informal way but it could be made precise using the notion of Baire generic. The set of Riemannian manifolds of a fixed dimension naturally can be made into a metric space by Nash embedding into a large Euclidean space $E$. A metric on Riemannian manifolds can then be defined by taking the smallest possible Hausdorff distance between any embeddings $M_{1},M_{2}$ in $E$. Alternatively one could fix a topological manifold $M$ and define a metric between Riemannian metrics on $M$.
@@
For a generic $d\geq 2$ dimensional Riemannian manifold with or without boundary, the wave front $W_{t}(P)$ becomes dense for all points $P\in M$.
@@ 4. Surfaces
We confirm here that one can have dense wave fronts even in very simple situations. Despite the fact that having a dense wave front displays a mixing property, it can even happen in the case when the geodesic flow is integrable. We look first at the flat 2-torus $M=\mathbb{R}^{2}/\mathbb{Z}^{2}$. It is a $2$-dimensional compact Riemannian manifold with Euclidean flat metric, induced from the plane.
@@
We look at the wave front on the universal cover $\mathbb{R}^{2}$ of $\mathbb{T}^{2}$, where it is a round circle of radius $t$. We can assume without loss of generality that the point $P$ is at the origin. The wave front can be seen as a union of two half circles which both are graphs. The first is the graph of $f(x)=\sqrt{t^{2}-x^{2}}$. The second is the graph of $g(x)=-\sqrt{t^{2}-x^{2}}$. We forget about the graph of $g$ and even on the graph of $f$, we concentrate only on a small rectangle $R$ of the circle which is chosen to be so wide that its height is $1$. For given $t>0$, define the region $R=[a,b]\times[f(a),f(b)]$ with $[a,b]=[-2\cdot\sqrt{t},-\sqrt{2t}]$. The graph of $W_{t}(p)$ intersected with $R$ crosses $R$ monotonically from $(a,f(a))$ to $(b,f(b))$. Since the graph of $f$ is concave down, the slope of $f$ is in the interval $I=[\sqrt{2},2]/\sqrt{1-t/4}$ which is for $t>36/5$ smaller than $3/\sqrt{t}$. We also check using a Taylor expansion of $f$ that that $f(b)-f(a)=1+O(1/\sqrt{t})$. This implies that $|f(t+1)-f(t)|\leq 3/\sqrt{t}$ in that interval $I$. If $R$ is projected to the fundamental domain $[0,1]\times[0,1]$, it crisscrosses that square and is $3/\sqrt{t}$ dense. We have shown that even a tiny part $W_{t}(P)\cap R$ of the entire wave front has become $3/\sqrt{t}$ dense. ∎
@@
4.4.
We investigated the torus problem first using results from the Gauss-circle problem which tells how many lattice points we expect in the annulus $A_{t,t+h}$ between $W_{t+h}$ and $W_{t}$ in the universal cover. Already Gauss estimated the number of lattice points in this annulus to be close to its area which is $2\pi th$ with an error $|E(t)|\leq\sqrt{2}\cdot 2\pi t$. It is known for example that $E(t)\leq Ct^{2/3}$.  [^1] The big open problem in this context is whether $E(t)=O(t^{1/2+\epsilon}$ holds for every $\epsilon>0$. This means that the annulus $A_{t,t+h}$ has about $2\pi th+Ct^{2/3}h$ lattice points. For $h=t^{-1/2}$, this area grows like $2\pi\sqrt{t}$ with an error $O(t^{1/6})$. As we expect so many lattice points in $A_{t,h}$ and every of these lattice points is $h=1/\sqrt{t}$ close to the wave front $W_{t}$ we see that $W_{t}$ gets $1/\sqrt{t}$ close to the original point $P$. Changing the annulus to $A_{t+h,t+2h}$ shows that we have points in distance $[1/\sqrt{t},2/\sqrt{t}]$ from $P$ etc. The geometry of numbers approach could even shed better light on how fast the wave fronts fill out the torus. The proof using the Gauss-circle problem in higher dimensions would need even heavier machinery like Ehrhard theory.
@@
The simple constructive proof given above was found only later in the summer of 2024. It illustrates that sometimes, it is easier to prove a stronger result: the proof shows that already a small part of the wave front becomes dense. The reader might enjoy to modify the simple single variable calculus argument to show that also on a stretched torus $\mathbb{R}/(\alpha\mathbb{Z})\times\mathbb{R}/(\beta\mathbb{Z})$, the wave front becomes dense. The argument can also be adapted to wave fronts on higher dimensional flat tori $\mathbb{R}^{d}/\mathbb{T}^{d}$. Again, we just have to restrict to a small cuboid domain $R$, where the wave front can be described in the cover as a graph of a function $f(x_{1},\dots,x_{d-1})$ of $d-1$ variables. There is no doubt that in the case of a non-flat metric, the density of the wave front persists, but this has not been proven; we do not know the shape of the wave front in the universal cover. But the odds appear to be good as the wave front gets longer. In the case when the geodesic flow has positive Lyapunov exponents, the wave front length in the universal cover would grow exponentially and make a density result more likely. But we even don’t know how to prove this. In the case when the geodesic flow is hyperbolic like on a compact Riemannian manifold of strictly negative curvature, the density would follow.
@@ 5. Billiards
Billiards in convex tables were used by George Birkhoff [9] (in section 6 of his book) and his postdoc Poritski [29] to illustrate principles and theorems in dynamical systems theory, especially in the context of the 3-body problem. Periodic points are critical points of a variational principle. The Birkhoff fixed point theorem even can establish the existence of many periodic points if the billiard map. For modern texts on billiards, see [34, 31]. A plethora of open problems persist: one does not know for example whether every triangular billiard has a periodic orbit. No smooth convex billiard exhibiting chaos (positive Lyapunov exponents on a set of positive measure in the phase space) is known. Candidates are tables like $x^{4}+y^{4}\leq 1$. Even one of the first problem of Poritski [29], asking whether every integrable smooth convex billiard is an ellipse, remains unsettled.
@@
There are Wiederkehr situations also for billiards, but it appears to be more restricted. If we start a wave front at the center $P$ of a circular billiard table, then the wave front will periodically return to the point $P$. But already for the same circle, we expect the wave front $W_{t}(P)$ to become dense, if the starting point $P$ is not at the center. We are not aware even whether the density of the wave front in the simple duck problem of the circle has been shown. We can not use the method used here. In general, we expect:
@@
[…]
@@
We were even unable to find references which ask such wave density questions. But we suspect that it must have been on the mind of many researchers in the field. Some numerical simulations of wave front beams have been done in [6, 5, 7]. These animations show the beauty and complexity which can happen with wave fronts, even for simple situations like a circle or an ellipse.
@@ 6. Polyhedra
The starting point of this investigation was a numerical simulation of the wave front on the unit cube. The second author has made movies of wave fronts on the cube since 2011 for multi-variable courses to illustrate how quickly simple open questions can appear in mathematics. A wave front on the torus can be seen on [25]. The problem of finding the smallest connection between two points of a cube or to find the path of a billiard shot in a rectangle reaching from a point to one of the corners by reflecting a given number of times on walls can all be solved by unwrapping the situation. This produces nice problems in problems related to distances in multi-variable calculus courses, like finding the shortest connection between two points on a cube.
@@
We were fascinated in the cube case by the fact that the wave fronts become more and more complicated and that the wave fronts $W_{t}(P)$ even do lose connectivity. Passing through a corner can cut the wave front because the geodesic flow is no more continuous there. In the universal cover of the cube, once several lattice points have been crossed, the wave front already has had a chance of been hacked into pieces. Computer experiments indicated that the wave front becomes dense even if we start at a corner. There is no Wiederkehr phenomenon also there. The biggest surprise was maybe that the lack of smoothness of the exponential map could leads to situations on polyhedra, where wave fronts become disconnected and complicated already after a relatively short time.
@@
The following result again follows from the corresponding result of the torus. The situation is tougher however. In the square case we dealt with a symmetry group that was the Klein group $G=(a,b,a^{2}=b^{2}=1,ab=ba)$ with 4 elements, which is a commutative group. In the cube case, we deal with the cube group a subgroup of $SO(3)$ of order of $24$. Every of the 6 faces of the cube comes in 4 possible rotated orders. If we draw a line $\gamma(t)$ connecting point $P$ with a point $Q$ in a square $X$, then depending on the crossing of the borders we attach a group element to this segment. There is so a rule which assigns to each point $Q$ of the plane an orientation and a point on the surface of the cube. We can model the situation as a 4:1 cover of the open flat manifold in which the 6 corners have been taken away. Imagine that each of the 6 open squares comes in 4 copies. Now these 24 squares can be glued together away from the corners to produce a smooth flat 2 manifold. If we look at the evolution of the wave front $W_{t}\cap Q$ in a face and denote by $W_{t}^{k}$ the by $k\pi/2$ turned front, then we can look at $W_{t}^{0}\cup W_{t}^{1}\cup W_{t}^{2}\cup W_{t}^{3}$ over each square. This union is now the union of 4 wave fronts on the torus. There is now a partition of $S_{t}$ in the universal cover into 4 sets $S_{t}^{k}$ which become increasingly dense each such that $\exp_{t}(S_{t}^{0})$ is the actual wave front $W_{t}(P)$ on the cube. By symmetry if one of the cases $\exp_{t}(S_{t}^{k})$ becomes dense, then all 4 do. Since the union does become dense by the theorem on the torus, also $W_{t}(P)$ does. We have shown:

### v2

@@ 1. Motivating questions
These pictures illustrate some rather general wave front problems in Riemannian geometry. One can ask for which Riemannian manifolds with or without boundary or on which polytopes do the wave front become dense for all initial points, like in the mirrored room example? In which cases does the wave front become dense for some points and not become dense for others like in the duck example case? In which cases does it not become dense for all initial points, like in the tsunami case?
@@
We explore this theme here in the torus, square and cube surface case, where the metric is flat and where the question can be analyzed with elementary tools only. While the subject could reach far into differential geometry, all we really need here is some single variable calculus.
Figure 1. What happens with a light wave front in a room of mirrors, a tsunami on a perfect sphere or a cube, and the wave front produced by a duck in a circular pond?
@@ 2. Background
The third framework considers piecewise smooth Riemannian polytops, a concept that was historically challenging to define [33, 29]. This difficulty has led either to restricting the analysis to convex polytopes [21, 41] or then to use a broader definition, where a polytop $M$ is viewed as the geometric realization of a finite abstract simplicial complex $G$, a finite set of non-empty sets closed under the operation of taking non-empty subsets [19]. A piecewise smooth Riemannian polytop is such a polytop that is equipped with a Riemannian metric $g$ that is smooth except on a subset of co-dimension $2$. Inductively, this means $M$ is a smooth manifold (with or without boundary) for which the tensor $g$ is defined on an open subset $U=M\setminus K$, where $K$ is a piecewise smooth Riemannian polytope of co-dimension $2$. The geodesic flow is defined for geodesic paths avoiding $K$. One of the simplest examples is the square table, a $2$-manifold with circular boundary and a metric that is discontinuous at its four corners. For a $3$-dimensional example, consider a billiard in a solid cube. Here, the Riemannian metric tensor $g$ may be discontinuous along the edges and vertices that form a 1-dimensional network.
@@ 3. Wave fronts
  • 2. Convex planar flat billiards. Now, $W_{t}$ remains a continuous curve. It remains piecewise smooth. Differentiability can fail at caustics or at the boundary.
  • 3. Piecewise flat convex polyhedra. In this case, $W_{t}$ in general can already become disconnected.
Figure 2. In a Riemannian d-manifold $M$, the wave front $W_{t}(P)$ of a point $P$ for which a tangent space $T_{P}M$ exists, is the image $\exp_{P}(S_{t})$ of the geodesic sphere $S_{t}=\{|x|=t\}\subset T_{P}M\sim\mathbb{R}^{d}$ of radius $t$.
@@
While for a smooth surface $M$ without boundary or a billiard in a convex region, the wave fronts $W_{t}$ are piecewise smooth and in particular continuous, this is already different on a closed polyhedral surface. The wave front there can become disconnected after some time. This happens in particular for the cube surface, the surface bounding the unit cube. Numerical experiments still indicate us also for more general non-flat compact surfaces, that in general, the wave front becomes dense, meaning that it fills out the surface more and more. An unresolved case is already the dodecahedron.
$3.$
Here is the definition of “dense wave front” for a general compact Riemannian manifold $M$ with or without boundary. Given $r>0$, the open ball $B_{r}(Q)$ of a point $Q\in M$ is the set of points $P$ for which there is a geodesic from $P$ to $Q$ that has geodesic length smaller than $r$.
@@
Since the (d1)(d-1)-dimensional volume of the co-dimension-one wave front $W_{t}$ is expected to grow and the manifold $M$ is finite, we expect it to fill up $M$ more and more. The term “generic” is here used in an informal way but it could be made precise using the notion of Baire generic. The set of Riemannian manifolds of a fixed dimension naturally can be made into a metric space by Nash embedding them into a large Euclidean space $E$. A metric on Riemannian manifolds can then be defined by taking the smallest possible Hausdorff distance between any embeddings $M_{1},M_{2}$ in $E$. Alternatively one could fix a topological manifold $M$ and define a distance between Riemannian metrics on $M$ and use this topology on metrics to define Baire genericity.
@@
For a generic $d\geq 2$-dimensional Riemannian manifold with or without boundary, the wave front $W_{t}(P)$ becomes dense for all points $P\in M$.
@@ 4. Surfaces
We confirm here that one can have dense wave fronts even in very simple situations. Despite the fact that having a dense wave front displays some sort of mixing property. This can even happen in the case when the geodesic flow is integrable. We look first at the flat 2-torus $M=\mathbb{R}^{2}/\mathbb{Z}^{2}$. It is a $2$-dimensional compact Riemannian manifold with Euclidean flat metric, induced from the plane.
@@
We look at the wave front on the universal cover $\mathbb{R}^{2}$ of $\mathbb{T}^{2}$, where it is a round circle of radius $t$. We can assume without loss of generality that the point $P$ is at the origin. The wave front can be seen as a union of two half circles which both are graphs. The first is the graph of $f(x)=\sqrt{t^{2}-x^{2}}$. The second is the graph of $g(x)=-\sqrt{t^{2}-x^{2}}$. We forget about the and even on the full graph of $f$, we concentrate only on a small rectangle $R$ of the circle which is chosen to be so wide that its height is $1$. For given $t>0$, define the region $R=[a,b]\times[f(a),f(b)]$ with $a,b]=[-2\cdot\sqrt{t},-\sqrt{2t}]$. The graph of $W_{t}(p)$ intersected with $R$ crosses $R$ monotonically from $(a,f(a))$ to $(b,f(b))$. Since the graph of $f$ is concave down, the slope of $f$ is in the interval $I=[\sqrt{2},2]/\sqrt{1-t/4}$ which is for $t>36/5$ smaller than $3/\sqrt{t}$. We also check using a Taylor expansion of $f$ that that $f(b)-f(a)=1+O(1/\sqrt{t})$. This implies that $|f(t+1)-f(t)|\leq 3/\sqrt{t}$ in that interval $I$. If $R$ is projected to the fundamental domain $0,1]\times[0,1]$, it crisscrosses that square and is $3/\sqrt{t}$ dense. We have shown that even a tiny part $W_{t}(P)\cap R$ of the entire wave front has become $3/\sqrt{t}$ dense. ∎
@@
$4.$
We investigated the torus problem first using results from the Gauss-circle problem which tells how many lattice points we expect in the annulus $A_{t,t+h}$ between $W_{t+h}$ and $W_{t}$ in the universal cover. Already Gauss estimated the number of lattice points in this annulus to be close to its area $2\pi th$ and estimated an error $|E(t)|\leq\sqrt{2}\cdot 2\pi t$. It is known for example that $E(t)\leq Ct^{2/3}$.  [^1] The big open problem in this context is whether $E(t)=O(t^{1/2+\epsilon}$ holds for every $\epsilon>0$. This means that the annulus $A_{t,t+h}$ has about $2\pi th+Ct^{2/3}h$ lattice points. For $h=t^{-1/2}$, this area grows like $2\pi\sqrt{t}$ with an error $O(t^{1/6})$. As we expect so many lattice points in $A_{t,h}$ and every of these lattice points is $h=1/\sqrt{t}$ close to the wave front $W_{t}$ we see that $W_{t}$ gets $1/\sqrt{t}$ close to the original point $P$. Changing the annulus to $A_{t+h,t+2h}$ shows that we have points in distance $1/\sqrt{t},2/\sqrt{t}]$ from $P$. The geometry of numbers approach could even shed better light on how fast the wave fronts fill out the torus. The proof using the Gauss-circle problem in higher dimensions would need heavier mathematical machinery like Ehrhard theory.
@@
The simple constructive proof given above was found only later in the summer of 2024. It illustrates that sometimes, it is easier to prove a stronger result: the proof shows that already a small part of the wave front becomes dense. The reader might enjoy to modify the simple single variable calculus argument to show that also on a stretched torus $\mathbb{R}/(\alpha\mathbb{Z})\times\mathbb{R}/(\beta\mathbb{Z})$, the wave front becomes dense. The argument can also be adapted to wave fronts on higher dimensional flat tori $\mathbb{R}^{d}/\mathbb{T}^{d}$. Again, we just have to restrict to a small cuboid domain $R$, where the wave front can be described in the cover as a graph of a function $f(x_{1},\dots,x_{d-1})$ of d1d-1 variables. There is no doubt that in the case of a non-flat metric, the density of the wave front persists, but this has not been proven; we do not know the shape of the wave front in the universal cover. But it is likely as the wave front gets longer. In the case when the geodesic flow has positive Lyapunov exponents, the wave front length in the universal cover would grow exponentially and make a density result more likely. But we even don’t know how to prove this. In the case when the geodesic flow is hyperbolic, like on a compact Riemannian manifold of strictly negative curvature, the density would follow.
@@ 5. Billiards
Billiards in convex tables were used by George Birkhoff [10] (in section 6 of his book) and his postdoc Hillel Poritsky [32] to illustrate principles and theorems in dynamical systems theory, especially in the context of the 3-body problem. Periodic points are critical points of a variational principle. The Birkhoff fixed point theorem even can establish the existence of many periodic points if the billiard map. For modern texts on billiards, see [37, 34]. A plethora of open problems persist: one does not know for example whether every triangular billiard has a periodic orbit. No smooth convex billiard exhibiting chaos (positive Lyapunov exponents on a set of positive measure in the phase space) is known. Candidates are tables like $x^{4}+y^{4}\leq 1$. Even one of the first problem of Poritski [32], asking whether every integrable smooth convex billiard is an ellipse, remains unsettled in its full form despite partial progress [9].
@@
There are Wiederkehr situations also for billiards, but it appears to be more restricted. If we start a wave front at the center $P$ of a circular billiard table, then the wave front will periodically return to the point $P$. But already for the same circle, we expect the wave front $W_{t}(P)$ to become dense, if the starting point $P$ is not at the center. We are not aware even whether the density of the wave front in the simple duck problem of the circle has been shown. We can not use the method used here for the torus. In general, we expect:
@@
By unwrapping the table we get to a universal cover $\mathbb{R}^{2}$ of $Q=[0,1]\times[0,1]$. The circle $S_{t}(P)$ centered at $P$ gets mapped into $K_{t}(P)=\pi(S_{t}(P))$, where $\pi$ is the map which first maps $(x,y)$ to $(x,y)\;{\rm mod}\;2$, then applies the map $(x,y)=(q(x),q(y))$, where $q$ is the tent map $q(x)=x$, if $x\leq 1$ and q(x)=2xq(x)=2-x if $x\in[1,2]$. In short, the piecewise smooth exponential map
$\pi(x,y)=(1-|(x\;{\rm mod}\;2)-1|,1-|(x\;{\rm mod}\;2)-1|)$
@@
We were even unable to find references which ask such wave density questions. But we suspect that it must have been on the mind of many researchers in the field. Some numerical simulations of wave front beams have been done in [6, 5, 7]. These animations show the beauty and complexity which can happen with wave fronts, even for simple situations like a circle or an ellipse. Related to wave fronts are caustics and the “Jacobi last theorem” type question of how many cusps such a caustic has. See [12].
@@ 6. Polyhedra
The starting point of this investigation was a numerical simulation of the wave front on the unit cube. The second author has made movies of wave fronts on the cube since 2011 for multi-variable courses to illustrate how quickly simple open questions can appear in mathematics. A wave front on the torus can be seen on [28]. The problem of finding the smallest connection between two points of a cube or to find the path of a billiard shot in a rectangle reaching from a point to one of the corners by reflecting a given number of times on walls can all be solved by unwrapping the situation. This produces nice puzzles in the context of distance questions in multi-variable calculus courses, like finding the shortest connection between two points on a cube or the length of a shortest billiard path in a rectangle that is required to bounce a pre-required number of times on each wall.
@@
We were fascinated to see in the cube case that the wave fronts become more and more complicated and that the wave fronts $W_{t}(P)$ even do lose connectivity. Passing through a corner can cut the wave front because the geodesic flow is no more continuous there. In the universal cover of the cube, once several lattice points have been crossed, the wave front already has had a chance of been hacked into pieces. Computer experiments indicated that the wave front becomes dense even if we start at a corner. There is no Wiederkehr phenomenon also there. The biggest surprise was maybe that the lack of smoothness of the exponential map could leads to situations on polyhedra, where wave fronts become disconnected and complicated already after a relatively short time.
@@
The following result again follows from the corresponding result of the torus. The situation is tougher however. In the square case we dealt with a symmetry group that was the Klein group $G=(a,b,a^{2}=b^{2}=1,ab=ba)$ with 4 elements, which is a commutative group. In the cube case, we deal with the cube group a finite subgroup of $SO(3)$ of order of $24$. Every of the 6 faces of the cube comes in 4 possible rotated orders. If we draw a line $\gamma(t)$ connecting point $P$ with a point $Q$ in a square $X$, then depending on the crossing of the borders we attach a group element to this segment. There is so a rule which assigns to each point $Q$ of the plane an orientation and a point on the surface of the cube. We can model the situation as a 4:1 cover of the open flat manifold in which the 6 corners have been taken away. Imagine that each of the 6 open squares comes in 4 copies. Now these 24 squares can be glued together away from the corners to produce a smooth flat 2 manifold. If we look at the evolution of the wave front $W_{t}\cap Q$ in a face and denote by $W_{t}^{k}$ the by $k\pi/2$ turned front, then we can look at $W_{t}^{0}\cup W_{t}^{1}\cup W_{t}^{2}\cup W_{t}^{3}$ over each square. This union is now the union of 4 wave fronts on the torus. There is now a partition of $S_{t}$ in the universal cover into 4 sets $S_{t}^{k}$ which become increasingly dense each such that $\exp_{t}(S_{t}^{0})$ is the actual wave front $W_{t}(P)$ on the cube. By symmetry if one of the cases $\exp_{t}(S_{t}^{k})$ becomes dense, then all 4 do. Since the union does become dense by the theorem on the torus, also $W_{t}(P)$ does. We have shown: